# FreeGraftor: Training-Free Cross-Image Feature Grafting for Subject-Driven Text-to-Image Generation

### v2

@@ Abstract.
Subject-driven image generation aims to synthesize novel scenes that faithfully preserve subject identity from reference images while adhering to textual guidance, yet existing methods struggle with a critical trade-off between fidelity and efficiency. Tuning-based approaches rely on time-consuming and resource-intensive subject-specific optimization, while zero-shot methods fail to maintain adequate subject consistency. In this work, we propose FreeGraftor, a training-free framework that addresses these limitations through cross-image feature grafting. Specifically, FreeGraftor employs semantic matching and position-constrained attention fusion to transfer visual details from reference subjects to the generated image. Additionally, our framework incorporates a novel noise initialization strategy to preserve geometry priors of reference subjects for robust feature matching. Extensive qualitative and quantitative experiments demonstrate that our method enables precise subject identity transfer while maintaining text-aligned scene synthesis. Without requiring model fine-tuning or additional training, FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both subject fidelity and text alignment. Furthermore, our framework can seamlessly extend to multi-subject generation, making it practical for real-world deployment. Our code is available at https://github.com/Nihukat/FreeGraftor.
copyright: noneconference: ; ;
@@
1. Introduction
The democratization of text-to-image generation has ignited widespread interest in personalized content creation, where users seek to synthesize novel scenes containing specific visual concepts—be it a cherished pet, a unique artifact, or a fictional character. While diffusion models(Nichol et al., 2021; Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2023) excel at generating diverse imagery from textual prompts, adapting them to preserve subject identity without costly per-subject optimization remains a critical challenge.
Current solutions for subject-driven generation(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) face a critical dilemma between fidelity and efficiency. Tuning-based methods, such as DreamBooth(Ruiz et al., 2023) and Textual Inversion(Gal et al., 2022), capture subject identity by fine-tuning models or learning embeddings on user-provided subject datasets. These methods rely on subject-specific optimization to achieve high subject consistency, which are resource-intensive and time-consuming, prone to catastrophic forgetting, and struggle with multi-subject composition. Zero-shot methods, such as IP-Adapter(Ye et al., 2023) and BLIP-Diffusion(Li et al., 2023), extract subject identities from reference images via additional trainable modules (e.g., image encoders or adapters) and integrate them into the text-to-image generation process. While these methods directly generate images containing reference subjects, they often fail to retain fine visual details, especially for novel concepts, as the limited parameter capacity of auxiliary modules fundamentally restricts their ability to reconstruct high-fidelity visual details.
It is worth noting that pre-trained text-to-image models inherently possess powerful visual representation capabilities, as evidenced by their ability to synthesize intricate textures and structures from textual prompts. However, current approaches fail to fully exploit this potential. Tuning-based methods modify the image generation distribution through fine-tuning, while zero-shot approaches offload visual representation learning to auxiliary modules. In contrast, directly tapping into the base model’s own feature space for cross-image knowledge transfer offers a promising way to bridge the gap between fidelity and efficiency. Recent advances in appearance transfer(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) lend credence to this idea, demonstrating that diffusion models can transplant visual attributes such as style and texture between images through semantic feature matching. For instance, Eye-for-an-eye(Go et al., 2024) successfully transfers localized appearances by establishing cross-image correspondences in U-Net’s feature space. By computing pixel-wise similarity (detailed in Sec. 3.2), we demonstrate that such cross-image semantic correspondences can also be identified in the feature space of modern MM-DiT-based(Esser et al., 2024) models (e.g., FLUX.1(Labs, 2024)). As illustrated in Figure 2, two pixels that are similar in FLUX.1’s feature space are also semantically corresponding.
Building on this insight, we propose FreeGraftor, a training-free framework that resolves the fidelity-efficiency dilemma in subject-driven generation via cross-image feature grafting. Our approach extracts visual features from reference images using rectified flow inversion(Rout et al., 2024; Wang et al., 2024b; Deng et al., 2024) and injects them into the generation process. Specifically, we introduce a novel cross-image feature grafting mechanism that transfers appearance details through semantic matching and position-constrained attention fusion in the feature space of text-to-image diffusion models. To bind reference patches to their corresponding generated patches, we implement a position embedding replication strategy, enabling semantically matched patches to share positional information during attention computation. Furthermore, we identify that proper noise initialization is crucial for robust semantic matching and structural preservation of reference subjects. Thus, we propose a structure-consistent initialization strategy: first constructing a collage containing the reference subject, then inverting it as the initialization noise for generation. Our method achieves pixel-level detail preservation (e.g., text, patterns) with fast inference and low computational overhead, requiring no additional training or optimization.
@@ 2. Related Work
2.1. Multimodal-Diffusion Transformer
Text-to-image diffusion models have achieved remarkable success in generating high-quality and diverse images, exemplified by DALL-E 2(Ramesh et al., 2022), Imagen(Saharia et al., 2022), GLIDE(Nichol et al., 2021), and eDiff-I(Balaji et al., 2022). Early popular open-source models, such as Stable Diffusion(Rombach et al., 2022) and SDXL(Podell et al., 2023), are based on U-Net(Ronneberger et al., 2015) architecture and utilize cross-attention between image and text to introduce textual conditions. In contrast, Diffusion Transformer (DiT)(Peebles and Xie, 2023) replaces the U-Net backbone with a Transformer(Vaswani et al., 2017), demonstrating superior scalability and performance. Recent advanced open-source models, including Stable Diffusion 3(Esser et al., 2024) and FLUX.1(Labs, 2024), follow this design and project text and image tokens into a unified space to compute joint attention, thereby better integrating complex relationships between text and image modalities. This architecture is termed Multimodal-Diffusion Transformer (MM-DiT). In this work, we implement our method on FLUX.1, where joint attention is computed as:
@@ 2.1. Multimodal-Diffusion Transformer
Here, [;][;][ ; ] denotes concatenation, $Q$, $K$, $V$ represent queries, keys, values, $PE$ is position embedding, $d$ denotes the dimension of the input features, $A$ denotes attention scores, and $H_{\text{out}}$ is the output of the attention layer.
@@ 2.2. Subject-Driven Generation
Subject-driven generation(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) aims to synthesize novel scenes containing user-specified subjects. Existing approaches fall into two categories: tuning-based and zero-shot. Tuning-based methods(Dong et al., 2022; Ruiz et al., 2023; Gal et al., 2022; Voynov et al., 2023; Alaluf et al., 2023; Liu et al., 2023; Kumari et al., 2023; Gu et al., 2023) achieve subject-driven generation by customizing models, typically requiring per-subject fine-tuning(Ruiz et al., 2023; Kumari et al., 2023; Gu et al., 2023) or iterative learning of unique embeddings(Gal et al., 2022; Voynov et al., 2023; Alaluf et al., 2023; Liu et al., 2023). While subject-specific tuning effectively captures visual details of the reference subject, it is resource-intensive and time-consuming, risks overfitting and catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods(Gal et al., 2023; Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Arar et al., 2023; Hua et al., 2023; Li et al., 2023; Ye et al., 2023; Ruiz et al., 2024; Wang et al., 2024a; Xiao et al., 2024) introduce additional learnable modules to process extra image inputs, seamlessly integrating visual information of the reference subject into the generation process. These methods enable personalization at inference time but sacrifice fine visual details, especially for novel concepts. Furthermore, they require large-scale retraining when switching to new base models. Moreover, both categories can degrade the performance of base models, resulting in reduced quality of generated images.
Beyond these categories, several training-free approaches(Ding et al., 2024; Shin et al., 2024) aim to harness the inherent capabilities of pre-trained text-to-image models for plug-and-play subject-driven generation. For example, FreeCustom(Ding et al., 2024) injects reference features by concatenating key and value pairs of reference and generated images in attention layers, emphasizing visual characteristics of the reference subject through attention score editing. However, this approach relies on layout alignment across multiple reference images and suffers from visual distortion due to coarse attention manipulation. DiptychPrompting(Shin et al., 2024) creatively reframes the subject-driven generation task as a diptych inpainting process, utilizing the visual context processing capabilities of pre-trained models to achieve high subject consistency. Unfortunately, pixel-space diptych construction incurs significant memory overhead, limiting its applicability to high-resolution images (e.g., 1024×1024 or larger). Additionally, this paradigm does not support generation with multiple reference subjects. In contrast, our work identifies cross-image semantic correspondences in MM-DiT’s feature space and binds matched patches via position embedding manipulation, enabling efficient subject feature injection with minimal computational cost.
@@ 2.3. Appearance Transfer
Appearance transfer methods(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) aim to combine the appearance of one image with the structure of another. Cross-Image(Alaluf et al., 2024) inject key-value pairs across images to transfer appearance from reference to target positions in the generated image. Dragon-diffusion(Mou et al., 2023) and DiffEditor(Mou et al., 2024) integrate noise guidance with key-value injection. However, these methods struggle to preserve structure because attention scores do not necessarily reflect semantic correspondences between reference and target images, potentially leading to erroneous matches. DIFT(Tang et al., 2023) identifies fine-grained semantic correspondences in U-Net feature space by computing pixel-pair similarities. Building on this, Eye-for-an-Eye(Go et al., 2024) achieves appearance transfer by directly replacing features of corresponding pixels. Nevertheless, these U-Net-based approaches are incompatible with modern MM-DiT architectures. Moreover, these methods focus on style transfer rather than identity preservation, which will cause geometric inconsistencies (e.g., shape, limb proportions) between reference and generated subjects if directly applied to subject-driven generation. In this work, we explore semantic matching and fine-grained feature transfer in MM-DiT while employing a noise initialization strategy to preserve the geometric structure of reference subjects.
Figure 3. Overview of FreeGraftor. First, we construct a collage based on the given text prompt and the reference subject image (Stage 1). Next, we invert this collage and record its diffusion trajectory (Stage 2). Finally, using the inverted noise as the initial latent representation, our FreeGraftor synthesizes the output image through iterative denoising. During this process, the Semantic-aware Feature Grafting (SAFG) module integrates features from the collage to ensure alignment with the reference subject (Stage 3).
Figure 4. Illustration of the Proposed Semantic-aware Feature Grafting (SAFG) Module. For each patch in the reference image, it first establishes semantic correspondence in the generated image via feature matching. The module then (1) concatenates the key and value pairs of the reference patch with those of the corresponding generated patch, while (2) copying and applying the position embedding from the generated patch to the reference patch’s key. This mechanism enables effective positional information sharing between corresponding patches while maintaining semantic alignment.
@@ 3.1. Overview
Our framework comprises three stages: collage construction, inversion, and generation (illustrated in Figure 3). First, we construct a collage using the given reference image and text prompt. Next, we invert the collage to obtain initialization noise for the generation phase while extracting reference features. Finally, these features are injected into the generation process to synthesize the final image containing the reference subject within the text-described scene.
3.2. Semantic-Aware Feature Grafting
Inspired by appearance transfer techniques(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024), we leverage the inherent capabilities of text-to-image diffusion models to transfer visual characteristics of reference subjects to corresponding regions in generated images without incurring additional training costs. To achieve this, we propose the Semantic-Aware Feature Grafting (SAFG) module, which first establishes semantic correspondences between reference and generated patches via semantic matching, then fuses features through position-constrained attention fusion (Figure 4).
3.2.1. Semantic Matching
3.2.2. Position-Constrained Attention Fusion
After obtaining the filtered reference mask $M^{\text{ref}}$, we transfer features from the retained reference patches. A naive approach is to directly replace the features of the corresponding patches. However, this may introduce artifacts along object boundaries or even cause image distortion (as shown in Figure 7). Unlike U-Net, which models positional relationships through convolutional operations, MM-DiT distinguishes patch proximity and orientation via position embeddings (PE). This enables us to align reference patches with their target positions in the generated image by editing position embeddings.
(9) $K^{\text{cat}}=[K^{\text{txt}};K^{\text{img,gen}};K_{M}^{\text{img,ref}}],$
$V^{\text{cat}}=[V^{\text{txt}};V^{\text{img,gen}};V_{M}^{\text{img,ref}}]$
For each retained reference patch $i$, we extract the corresponding position embedding $PE_{m(i)}^{\text{img, gen}}$ based on the matching relationship $m(i)$. The new position embedding is then formed by concatenating:
The revised attention computation is formulated as:
Here, $A^{+}$ represents the updated attention scores, and $H_{\text{out}}^{+}$ denotes the output of the attention layer. By integrating semantic matching and the sharing of position embeddings, our Semantic-Aware Feature Grafting module effectively binds features between matching patches in the reference and generated images, facilitating cross-image feature grafting.
3.3. Structure-Consistent Initialization
Attention-based feature migration can replicate the appearance of reference subjects (e.g., color, texture) but may not preserve structural details (e.g., shape, body proportions). Additionally, layouts generated from Gaussian noise exhibit significant randomness, leading to discrepancies in shape and size between generated subjects and those in the reference image, potentially causing semantic matching failures. To maintain structural integrity and enhance the robustness of semantic matching, we propose a structure-consistent initialization strategy. This involves creating a collage containing the reference subject and then employing an inversion technique to derive the initial noise required for the generation phase.
As illustrated in Figure 3, given a text prompt (e.g., ”a cat sitting on a chair wearing a birthday hat”), we first use a base text-to-image model to generate a template image that faithfully represents the described scene. To preserve the structure of the reference subject, we replace the corresponding subject in the template image with the reference subject. Specifically, a grounding model localizes the target subject in the template, a segmentation model extracts its mask, and an inpainting model removes the subject from the template. The reference subject is then cropped (using the same grounding and segmentation models), resized, and pasted into the erased region to form a collage. This collage replaces the original reference image as the new reference. Finally, we invert this collage to obtain the initial noise for the generation phase and record the diffusion trajectory during inversion to facilitate feature grafting.
To prevent the generated subject from overfitting the pose of the reference subject, we introduce a dynamic feature dropout strategy. Since structural layouts are predominantly determined during early diffusion steps(Balaji et al., 2022; Alaluf et al., 2024), we reduce reference feature injection in early steps while retaining late-step features to preserve visual details. At timestep $t$, for each reference patch, we apply dropout with probability $p=\omega t$, where $\omega\in[0,1]$ controls dropout intensity. The dropout mask is represented by a Bernoulli distribution:
(14) ${M^{\text{ref}}}^{\prime}=M^{\text{ref}}\odot M^{\text{drop}}$
This strategy preserves the reference subject’s structure while aligning it with the text-driven layout and maintaining pose flexibility. For multi-subject generation, we iteratively replace subjects in the template and invert the multi-subject collage, avoiding simultaneous processing of multiple reference images and eliminating computational overhead.
Figure 8. Generation results under different similarity threshold ($\tau$) and cycle-consistency threshold ($\delta$) settings. Higher $\tau$ and lower $\delta$ lead to more stringent filtering.
Figure 9. Generation results under different dropout intensity ($\omega$) settings. Higher $\omega$ improves text alignment but degrades subject consistency.
4. Experiments
4.1. Experimental Setup
4.1.1. Implementation Details
Our method is implemented on FLUX.1-dev. For collage construction, we use Grounding DINO(Liu et al., 2024) to localize subjects, SAM(Kirillov et al., 2023) to segment subjects, and LaMa(Suvorov et al., 2022) to erase the original subject in the template image. For rectified flow inversion, FireFlow(Deng et al., 2024) is employed as the inversion method. Template generation, inversion, and final generation each take 25 diffusion steps. During semantic matching, we set the similarity threshold $\tau=0.2$, cycle consistency threshold $\delta=1.5$, and dropout intensity $\omega=0.5$.
4.1.2. Baselines
We compare our method with previous subject-driven generation approaches, including zero-shot methods (IP-Adapter(Ye et al., 2023), MS-Diffusion(Wang et al., 2024a), OmniGen(Xiao et al., 2024)) and training-free methods (FreeCustom(Ding et al., 2024), DiptychPrompting(Shin et al., 2024)). Detailed descriptions of these methods are provided in Appendix B.
4.1.3. Datasets and Evaluation
For qualitative evaluation, we use reference images from DreamBench(Ruiz et al., 2023), CustomConcept101(Kumari et al., 2023), and Mix-of-Show(Gu et al., 2023) for subject-driven generation. For quantitative evaluation, we generate 30×25×4=3,000 images using 30 subjects and 25 prompt templates from DreamBench(Ruiz et al., 2023), with 4 random seeds per subject-prompt pair. We evaluate performance from two perspectives: text alignment and image alignment. For text alignment, we report results using CLIP(Radford et al., 2021) and ImageReward(Xu et al., 2024) metrics. For image alignment, we use GroundedSAM(Ren et al., 2024) to extract segments of reference and generated subjects, respectively, and compute similarity scores between these segments using CLIP(Radford et al., 2021) and DINOv2(Oquab et al., 2023).
4.2. Quantitative Results
Image-alignment Text-alignment
Method CLIP-I DINO CLIP-T ImageReward
FreeCustom 0.8308 0.6107 0.3246 0.6223
IP-Adapter 0.8920 0.7696 0.3048 0.7444
MS-Diffusion 0.9023 0.7977 0.3254 1.3405
OmniGen 0.9113 0.8167 0.3256 1.4926
DiptychPrompting 0.8924 0.7971 0.3291 1.5728
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
w/o Init 0.8529 0.7138 0.3236 1.6320
w/o Graft 0.8463 0.6572 0.3306 1.6434
w/o Match 0.9504 0.8974 0.3232 1.5258
w/ replace 0.8476 0.7010 0.3174 1.4551
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
As shown in Table 1, FreeGraftor demonstrates superior image alignment and text alignment compared to prior methods. The significant improvement in image alignment indicates our method effectively preserves the identity of reference subjects: structure-consistent initialization retains geometric priors, while feature grafting successfully transfers fine-grained visual details. The high text alignment stems from leveraging the base text-to-image model without fine-tuning or training auxiliary modules, thereby preserving its original capabilities. Furthermore, FreeGraftor employs a gentle attention fusion strategy to inject subject features, minimizing perturbations to vanilla attention computation.
4.3. Qualitative Results
4.3.1. Single-Subject Generation
Figure 5 presents the generation results for single-subject generation using FreeGraftor and other methods. FreeCustom and IP-Adapter fail to retain fine visual details of the reference subject effectively. While MS-Diffusion, OmniGen, and DiptychPrompting partially recover reference characteristics, they struggle with pixel-level fidelity. FreeGraftor achieves pixel-accurate detail preservation, such as logos on the can, numerals on the clock, and text on the teddy bear. For non-rigid subjects (e.g., dogs, plush toys), FreeGraftor generates natural pose variations aligned with text guidance, demonstrating the flexibility and robustness of our feature grafting technique.
4.3.2. Multi-Subject Generation
As IP-Adapter and DiptychPrompting do not support multi-subject generation, we compare FreeGraftor with FreeCustom, MS-Diffusion, and OmniGen. As shown in Figure 6, FreeCustom exhibits severe attribute confusion (e.g., facial structures of the cat and the dog, hairstyles of the man and the woman). MS-Diffusion and OmniGen show higher fidelity but still lack visual faithfulness for small objects (e.g., the cup). FreeGraftor preserves all reference details while enabling flexible text-guided control, highlighting its effectiveness in multi-subject generation.
4.4. Ablation Study
4.4.1. Component Ablation
To validate the effectiveness of each component in our method, we compare several ablation variants, as shown in Figure 7. (a) When structure-consistent initialization is removed, the generated images inherit the reference subject’s appearance (e.g., colors, patterns) but fail to maintain structural consistency (e.g., distorted body proportions in dogs). (b) Without feature grafting, the generated subjects lack identifiable correspondence with the reference subjects. (c) If semantic matching is omitted and reference keys and values are directly appended during attention computation, the generated subjects exhibit rigid poses and artifacts. (d) Replacing position-constrained attention fusion with direct feature replacement introduces incoherent edges and distortions. Quantitative results in Table 2 confirm that both structure-consistent initialization and feature grafting significantly improve image alignment, as they are critical for preserving subject identity. Skipping semantic matching or positional fusion degrades text alignment due to reduced image quality.
4.4.2. Hyperparameter Analysis
We conduct additional ablation experiments to analyze the impact of similarity threshold $\tau$ and cycle consistency threshold $\delta$ on semantic matching. Appropriate thresholds alleviate mismatches (e.g., unnatural expansion of the backpack), while overly strict thresholds ($\tau\uparrow$ or $\delta\downarrow$) discard critical visual details (e.g., missing badges on the backpack, Figure 8).
We further verify the effectiveness of dropout in mitigating structural overfitting. As shown in Figure 9, moderate dropout enhances pose flexibility while preserving subject identity, whereas excessive dropout intensities ($\omega>0.6$) erode identity retention. These findings demonstrate that through careful hyperparameter selection, our method achieves an optimal balance between pose flexibility and detail preservation.
5. Conclusion
We present FreeGraftor, a novel training-free framework for subject-driven generation that bridges the critical gap between subject fidelity and computational efficiency. By leveraging Semantic-Aware Feature Grafting, FreeGraftor seamlessly injects the appearance of reference subjects into generated scenes. Additionally, our Structure-Consistent Initialization strategy preserves reference geometry and enhances the robustness of semantic matching. Extensive experiments demonstrate that FreeGraftor achieves pixel-level detail preservation and flexible text-guided control without requiring any training or test-time optimization. Our framework can be seamlessly extended to multi-subject generation, faithfully retaining visual details of all reference subjects while incurring no additional computational overhead.
Appendix A Pseudo Code

### v3

@@ Abstract.
Subject-driven image generation aims to synthesize novel scenes that faithfully preserve subject identity from reference images while adhering to textual guidance. However, existing methods struggle with a critical trade-off between fidelity and efficiency. Tuning-based approaches rely on time-consuming and resource-intensive, subject-specific optimization, while zero-shot methods often fail to maintain adequate subject consistency. In this work, we propose FreeGraftor, a training-free framework that addresses these limitations through cross-image feature grafting. Specifically, FreeGraftor leverages semantic matching and position-constrained attention fusion to transfer visual details from reference subjects to the generated images. Additionally, our framework introduces a novel noise initialization strategy to preserve the geometry priors of reference subjects, facilitating robust feature matching. Extensive qualitative and quantitative experiments demonstrate that our method enables precise subject identity transfer while maintaining text-aligned scene synthesis. Without requiring model fine-tuning or additional training, FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both subject fidelity and text alignment. Furthermore, our framework can seamlessly extend to multi-subject generation, making it practical for real-world deployment. Our code is available at https://github.com/Nihukat/FreeGraftor.
subject-driven image generation, text-to-image synthesis, training-free framework
$c$s: Computing methodologies Image manipulation
1. Introduction
The democratization of text-to-image generation has ignited widespread interest in personalized content creation, where users seek to synthesize novel scenes containing specific visual concepts—be it a cherished pet, a unique artifact, or a fictional character. While diffusion models(Nichol et al., 2021; Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2023) excel at generating diverse imagery from textual prompts, adapting them to preserve subject identity without incurring costly per-subject optimization remains a critical challenge.
Current subject-driven generation methods(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) face a critical dilemma between fidelity and efficiency. Tuning-based methods (e.g., DreamBooth(Ruiz et al., 2023)) achieve high subject consistency by fine-tuning the model on reference images. However, this subject-specific optimization is resource-intensive, time-consuming, prone to catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods (e.g., IP-Adapter(Ye et al., 2023)) circumvent fine-tuning by using auxiliary modules (e.g., image encoders or adapters) to extract and inject subject identity information into the generation process. While efficient, these methods often fail to retain fine visual details, particularly for novel concepts, as the limited capacity of the auxiliary modules fundamentally restricts high-fidelity reconstruction.
Critically, pre-trained text-to-image models inherently possess powerful visual representation capabilities, as evidenced by their ability to synthesize intricate textures and structures directly from text. However, current approaches largely overlook this intrinsic potential: tuning-based methods shift the generation distribution via fine-tuning, while zero-shot methods delegate representation learning to additional modules. Instead, directly leveraging the base model’s own rich feature space for cross-image knowledge transfer presents a promising avenue for bridging the fidelity-efficiency gap. Recent advances in appearance transfer(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) support this direction, demonstrating that diffusion models can transplant visual attributes (e.g., style, texture) between images through semantic feature matching. For instance, Eye-for-an-eye (Go et al., 2024) successfully transfers localized appearances by establishing cross-image correspondences within the U-Net’s feature space.
Building upon this concept, we identify a key insight: such cross-image semantic correspondences are not limited to U-Net architectures but also exist robustly within the feature space of modern MM-DiT-based models (Esser et al., 2024) (e.g., FLUX.1 (Labs, 2024)). As demonstrated in Fig. 1, computing pixel-wise similarity (detailed in Sec. 3.2) reveals that semantically corresponding pixels exhibit high similarity in FLUX.1’s feature space.
@@ 1. Introduction
Leveraging this insight, we propose FreeGraftor, a training-free framework that resolves the fidelity-efficiency dilemma in subject-driven generation via cross-image feature grafting. Our approach extracts visual features from reference images using rectified flow inversion (Rout et al., 2024; Wang et al., 2024b; Deng et al., 2024) and injects them into the generation process. Specifically, we introduce a novel cross-image feature grafting mechanism that transfers appearance details through semantic matching and position-constrained attention fusion in the feature space of text-to-image diffusion models. To bind reference patches to their corresponding generated patches, we implement a position embedding replication strategy, enabling semantically matched patches to share positional information during attention computation. Furthermore, we identify that proper noise initialization is crucial for robust semantic matching and structural preservation of reference subjects. Thus, we propose a structure-consistent initialization strategy: first constructing a collage containing the reference subject, then inverting it as the initialization noise for generation. Our method achieves pixel-level detail preservation (e.g., text, patterns) with fast inference and low computational overhead, requiring no additional training or optimization.
@@ 2. Related Work
2.1. Text-to-Image Diffusion Models
Text-to-image diffusion models(Ramesh et al., 2022; Saharia et al., 2022; Nichol et al., 2021; Balaji et al., 2022) have achieved remarkable success in generating high-quality and diverse images. Early popular open-source models, such as Stable Diffusion(Rombach et al., 2022) and SDXL(Podell et al., 2023), are based on the U-Net(Ronneberger et al., 2015) architecture and utilize cross-attention between image and text to incorporate textual conditions. In contrast, Diffusion Transformer (DiT)(Peebles and Xie, 2023) replaces the U-Net backbone with a Transformer(Vaswani et al., 2017), demonstrating superior scalability and performance. Recent Multimodal-Diffusion Transformers (MM-DiTs), including Stable Diffusion 3(Esser et al., 2024) and FLUX.1(Labs, 2024), follow this design and project text and image tokens into a unified space to compute joint attention, thereby better integrating complex relationships between text and images. In this work, we implement our method on FLUX.1, where joint attention is computed as:
@@ 2.1. Text-to-Image Diffusion Models
Here, $[;]$ denotes concatenation, $Q$, $K$, $V$ represent queries, keys, values, $PE$ is position embedding, $d$ denotes the dimension of the input features, $A$ denotes attention scores, and $H_{\text{out}}$ is the output of the attention layer.
@@ 2.2. Subject-Driven Generation
Subject-driven generation(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) aims to synthesize novel scenes containing user-specified subjects. Early textual inversion methods(Gal et al., 2022; Dong et al., 2022; Voynov et al., 2023; Alaluf et al., 2023; Liu et al., 2023) learn new concepts by iteratively optimizing embeddings. Mainstream approaches fall into two categories: tuning-based and zero-shot. Tuning-based methods(Ruiz et al., 2023; Kumari et al., 2023; Gu et al., 2023) require per-subject fine-tuning to effectively capture visual details of the reference subject, which is resource-intensive and time-consuming, risks overfitting and catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods(Gal et al., 2023; Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Arar et al., 2023; Hua et al., 2023; Li et al., 2023; Ye et al., 2023; Ruiz et al., 2024; Wang et al., 2024a; Wu et al., 2025; Mou et al., 2025) introduce additional learnable modules to process extra image inputs, enabling personalization at inference time. Most of them rely on external encoders to compress images into embeddings, achieving only semantic-level alignment and suffering from insufficient visual consistency, particularly for novel concepts. UNO(Wu et al., 2025) and DreamO(Mou et al., 2025) leverage the VAE within base models to encode reference images for better subject consistency, yet they still struggle to retain fine-grained visual details (e.g., text and patterns). Furthermore, zero-shot methods require large-scale retraining when switching to new base models. Moreover, both categories can degrade the performance of base models, resulting in reduced quality of generated images.
Beyond these categories, several training-free methods(Ding et al., 2024; Shin et al., 2024; Kang et al., 2025) aim to harness the in-context learning potential of pre-trained text-to-image models for plug-and-play subject-driven generation. FreeCustom(Ding et al., 2024) concatenates the reference images and noise in the latent space, and enhances the generated image’s focus on reference subjects by implementing a weighted mask strategy within the self-attention mechanism. DiptychPrompting(Shin et al., 2024) extends this paradigm to MM-DiT-based models via a ControlNet(Zhang et al., 2023), and reframes the subject-driven generation task as image completion on a diptych through meticulous prompt design. LatentUnfold(Kang et al., 2025) dispenses with ControlNet dependencies and evolves the diptych into a grid to enhance subject consistency. These methods avoid resource-demanding training or fine-tuning. However, constrained by the limited in-context learning capacity of base models, relying solely on self-attention struggles to fully reconstruct every detail of the reference subject. Our method identifies cross-image semantic correspondences while binding matched patches, reinforcing contextual connections between reference and generated subjects, thereby ensuring consistency across all components of the subject.
@@ 2.3. Appearance Transfer
Appearance transfer methods(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) aim to combine the appearance of one image with the structure of another. Techniques like Cross-Image (Alaluf et al., 2024) primarily rely on injecting key-value pairs across images to transfer appearance from reference to target positions in the generated image. Dragon-diffusion(Mou et al., 2023) and DiffEditor(Mou et al., 2024) integrate noise guidance with key-value injection. However, these methods often struggle to preserve structure accurately, as attention scores do not reliably reflect semantic correspondences, which can lead to erroneous matches. Methods like DIFT (Tang et al., 2023) address this by identifying fine-grained correspondences through pixel-pair similarity in U-Net’s feature space. Building on this, Eye-for-an-Eye(Go et al., 2024) achieves appearance transfer by directly replacing features of corresponding pixels. Nevertheless, these methods focus on style transfer rather than identity preservation, which can cause geometric inconsistencies (e.g., shape, limb proportions) between reference and generated subjects. Besides, these U-Net-based approaches are incompatible with modern MM-DiT architectures. In contrast, our work explores semantic matching and fine-grained feature transfer within the MM-DiT architecture while employing a dedicated noise initialization strategy to preserve the geometric structure of reference subjects.
Figure 3. Overview of FreeGraftor. First, we construct a collage based on the given text prompt and the reference image (Stage 1). Next, we invert this collage and record its diffusion trajectory (Stage 2). Finally, using the inverted noise as the initial latent representation, FreeGraftor synthesizes the output image through iterative denoising. During this process, the Semantic-aware Feature Grafting (SAFG) module integrates features from the collage to ensure alignment with the reference subject (Stage 3).
@@ 3.1. Overview
Our framework comprises three stages: collage construction, inversion, and generation (illustrated in Fig. 3). First, we construct a collage using the given reference image and text prompt. Next, we invert the collage to obtain initialization noise for the generation phase while extracting reference features. Finally, these features are injected into the generation process to synthesize the final image that contains the reference subject within the text-described scene.